OpenStudy (anonymous):

If sin A = 2/3 where 0° < A < 90°, what is the value of sin 2A

OpenStudy (anonymous):

remember the double angle formula: \[\sin(2A)=2\sin A\cos A\]we already have the \(\sin A\) part and need the cosA. to get that, we use the identity \[\sin^2A+\cos^2A=1\] using this, we get two possible values for cosine -> one positive and one negative. Since the angle A is in the first quadrant, cosA must be positive and we take the positive value. then plug in the numbers and find sin 2A